### v1

# A Watermark for Low-entropy and Unbiased Generation in Large Language Models

Abstract
Recent advancements in large language models (LLMs) have highlighted the risk of misuse, raising concerns about accurately detecting LLM-generated content. A viable solution for the detection problem is to inject imperceptible identifiers into LLMs, known as watermarks. Previous work demonstrates that unbiased watermarks ensure unforgeability and preserve text quality by maintaining the expectation of the LLM output probability distribution. However, previous unbiased watermarking methods are impractical for local deployment because they rely on accesses to white-box LLMs and input prompts during detection. Moreover, these methods fail to provide statistical guarantees for the type II error of watermark detection. This study proposes the Sampling One Then Accepting (STA-1) method, an unbiased watermark that does not require access to LLMs nor prompts during detection and has statistical guarantees for the type II error. Moreover, we propose a novel tradeoff between watermark strength and text quality in unbiased watermarks. We show that in low-entropy scenarios, unbiased watermarks face a tradeoff between watermark strength and the risk of unsatisfactory outputs. Experimental results on low-entropy and high-entropy datasets demonstrate that STA-1 achieves text quality and watermark strength comparable to existing unbiased watermarks, with a low risk of unsatisfactory outputs. Implementation codes for this study are available online. [^1] https://github.com/djwei96/STA
1 Introduction
Large language models (LLMs) are large-scale deep learning models that can understand and generate natural languages by learning from a large amount of textual data. Typical generative LLMs, such as ChatGPT [20] and LLaMA [29], can answer questions, translate languages, and create codes with qualities comparable to humans. As LLMs can generate contents more efficiently at a lower cost compared to humans, the risk of LLMs being employed to generate biased, fake, or malicious contents is also increasing [19]. For example, LLMs may exhibit biased information against underrepresented groups of people [2, 7], create misinformation [21], and harm academic integrity [35]. To reduce the harm caused by LLMs, identifying LLM-generated content precisely and efficiently becomes a crucial issue [12].
Watermarks are identifiers imperceptible to humans but detectable by certain models [18]. In particular, in the era of LLMs, a watermarking method is a strategy to alter the logits or probability distribution during token generation by LLMs [12, 8], with the randomness controlled confidentially by LLM owners. Statistical hypotheses are constructed with the knowledge advantage for the randomness (as the alternative hypothesis) and without the knowledge advantage (as the null hypothesis). For example, previous work (denoted as KGW) [12] first divides the token set into a green and a red list at each generation step. Next, it increases the logits output by the pretrained LLM for green list tokens, leading the watermarked LLM to generate tokens in the green list with higher probabilities. The knowledge advantage in KGW is the randomness of the green-red-list partition at each step.
In practice, watermarks should have unforgeability against deciphering watermarking generation attacks [18]. A watermarking method demonstrates the ability against forgeries if it can hide the distinguishability between the original unwatermarked text and its watermarked counterpart [5, 18]. Thus, it is required that a watermarking method adjusts the probability distribution while maintaining the same expectation as the unwatermarked distribution [11, 15], defined as unbiased watermarks.
Hu et al. [11] find KGW is biased under the expectation maintaining requirement. In response, they propose an unbiased $\gamma$-reweight method that alters the probability distribution at each generation step. To detect watermarked texts, they use the log likelihood ratio (LLR) test by comparing the likelihood of the text produced by watermarked and unwatermarked white-box LLMs. Their knowledge advantage to construct statistical tests is the controlled randomness in $\gamma$-reweight. Wu et al. [33] propose Dipmark, an extension of $\gamma$-reweight with more general parameter settings.
Existing unbiased watermarking methods focus on the unbiased reweighting of the initial probability distribution [11, 33]. However, these methods face challenges for the following reasons. First, to accurately compute the likelihood ratio, methods employing the LLR test require the prompt as input and a white-box LLM even in watermark detection [8, 11]. The requirement of a white-box LLM is costly and impractical for local deployment. Also, relying on the LLR test makes watermarks unstable because changing the first token of the generated text can lead to huge deviations from the original likelihood value, due to the autoregressive mechanism of LLM inference [8]. Second, although both $\gamma$-reweight and Dipmark ensure the type I error of watermark detection, they fail to provide statistical guarantees for the type II error [11, 33]. That is, they can only guarantee that unwatermarked texts are not misclassified, but cannot guarantee that all watermarked texts are detectable.
In response to the above challenges, we propose the Sampling One Then Accepting (STA-1) method for watermarking LLMs. STA-1 traces back to the original KGW method where the token set is divided into a green and a red list at each generation step [12]. Instead of raising logits in the green list, STA-1 samples a token from the original probability distribution and accepts it if it is in the green list. If the sampled token is in the red list, it resamples another token and accept it. By counting the number of green list tokens, it employs the $z$-test for watermark detection, which naturally addresses the robustness issue encountered with the LLR test and eliminates the need for prompts and white-box LLMs in detection [12]. Meanwhile, we prove that STA-1 is unbiased and provides implicit statistical guarantees for the type II error. More interestingly, the bounds are linked to the Gini index of the probability distribution, which is a common metric in machine learning [3] compared to the proposed Spike entropy in previous work [12].
In this study, we also clarify the watermark strength and text quality tradeoff in unbiased watermarks. The KGW method faces a tradeoff between watermark strength and text quality, which means a higher detection power results in a lower text quality [12]. Previous work claims that unbiased watermarks can avoid this tradeoff given the preserved text quality by maintaining the expectation of probability distribution [11]. We challenge this claim by considering a simple low-entropy scenario, where we show that unbiased watermarks still face a tradeoff between watermark strength and text quality. However, the text quality is now related to the risk of unsatisfactory outputs. Specifically, unsatisfactory outputs in low-entropy scenarios represent that the watermarking method alters probability distribution such that high-probability tokens cannot be sampled at risk. We discuss the risk via the variance of the probability after altering, which is a common practice of risk-return analysis [27]. We prove that STA-1 is less risky than previous unbiased watermarks. Moreover, we propose STA-M, an extension of STA-1, by setting up a threshold for entropy in generation [16, 30] and sampling more times for high-entropy steps. Although STA-M is not unbiased theoretically, it allows higher watermark strength with small performance shifts empirically. Also, STA-M is robust against various watermarking attacks. We summarize our contributions as follows:
  • We propose STA-1, an unbiased watermarking method that is practical and has statistical guarantees on type II error of watermark detection. Moreover, we introduce STA-M, an extension of STA-1 that enhances watermark strength with low text quality shifts.
  • We clarify the watermark strength and text quality tradeoff in unbiased watermarks. In low-entropy scenarios, the text quality is related to the risk of unsatisfactory outputs. We show that STA-1 has a lower risk theoretically compared to other unbiased watermarks.
  • Experimental results on public low-entropy and high-entropy datasets empirically show that STA-1 achieves comparable performances against other unbiased watermarks and has a low risk of unsatisfactory outputs. Meanwhile, STA-M demonstrates high watermark strength in the low-entropy dataset and robustness against different watermarking attacks.
2 Preliminary
Notations. We follow notations in previous work [12, 11] to represent the generation task of LLMs. Let $P_{M}$ denote a pretrained LLM and $\mathcal{V}$ is the overall token (vocabulary) set. An example token set contains more than 50,000 tokens ($|\mathcal{V}|>50000$) [23]. For simplicity, we use Python-style notation for an ordered token sequence, where $x^{-m:n}=(x^{-m},x^{-m+1},\cdots,x^{n})$, $m$ and $n$ are integers. In a typical LLM generation task, an LLM receives a sequence of $N_{p}+1$ tokens $x^{-N_{p}:0}$, known as a prompt, and outputs a sequence of $T$ tokens $x^{1:T}$ step by step. At step $t$, the probability of each token in the token set $\mathcal{V}$ is given by the conditional distribution $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$. The LLM generation follows an autoregressive fashion, where the joint probability of the generated tokens are as follows
$P_{M}(x^{1:T}|x^{-N_{p}:(-1)})=\prod_{t=1}^{T}P_{M}(x^{t}|x^{-N_{p}:(t-1)}).$ (1)
When applying watermarking techniques, the LLM employs a private key $k$ to adjust the conditional distribution from $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ to $P_{M,w}(x^{t}|x^{-N_{p}:(t-1)};k)$, where $P_{M,w}$ indicates a watermarked model and the private key $k$ is randomly selected from a key space $K$ according to a known distribution $P_{K}(k)$. An unbiased watermark requires that the expectation of the watermarked distribution equals that of the original distribution [11, 18], defined as follows.
Definition 1 (Unbiased watermark). Given a prompt $x^{-N_{p}:0}$ and a known distribution $P_{K}(k)$ of the key $k$, a watermarking method is unbiased towards the original model $P_{M}$ if the watermarked model $P_{M,w}$ satisfies
$\mathbb{E}_{k\sim P_{K}(k)}\left[P_{M,w}(x^{t}|x^{-N_{p}:(t-1)};k)\right]=P_{M}(x^{t}|x^{-N_{p}:(t-1)}),$ (2)
for any prompt $x^{-N_{p}:0}\in\mathcal{V}^{N_{p}+1}$, any token $x^{t}\in\mathcal{V}$, and all generation steps $1\leq t\leq T$.
Previous distribution reweighting methods. Many existing watermarking methods fall into the category of distribution reweighting [11, 12, 16, 33]. Formally, a reweighting function $R_{k}:\mathcal{P}_{\mathcal{V}}\rightarrow\mathcal{P}_{\mathcal{V}}$ maps from $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ to $P_{M,w}(x^{t}|x^{-N_{p}:(t-1)};k)$, where $\mathcal{P}_{\mathcal{V}}$ denotes the probability distribution space over the vocabulary set $\mathcal{V}$. A reweighting method $R:K\times\mathcal{P}_{\mathcal{V}}\rightarrow\mathcal{P}_{\mathcal{V}}$ contains all realized reweighting functions $R_{k}$ among the key space $k\in K$. Following Definition 1, $R$ is an unbiased reweighting method if $\mathbb{E}_{k\sim P_{K}(k)}\left[R_{k}(P_{M})\right]=P_{M}$. Next, we introduce previous distribution reweighting methods [11, 12, 33] and refer readers to Appendix A for more details.
KGW [12]: The KGW method [12] randomly splits the vocabulary set $\mathcal{V}$ into a green list and a red list based on a uniformly distributed key $k$. The soft KGW method adds a predefined constant $\delta$ to the green list tokens’ logits while keeping the red list tokens’ logits fixed.
Dipmark [33] and $\gamma$-reweight [11]: Wu et al. [33] propose an unbiased watermarking method named Dipmark. Dipmark shuffles all probability masses $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ over the vocabulary set within the probability interval $[0,1]$ based on a key $k$. A hyperparameter $\alpha\in[0,0.5]$ partitions the interval $[0,1]$ into three segments: $[0,\alpha]$, $(\alpha,1-\alpha]$, and $(1-\alpha,1]$. Probabilities in the first segment are set to 0, those in the second remain constant, and those in the third are doubled. Dipmark becomes $\gamma$-reweight when $\alpha=0.5$.
3 A Simple Protocol for a Low-entropy Scenario
The low-entropy text refers to a relatively deterministic sequence in natural language. The entropy measures the uncertainty of the probability distribution $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ at a single generation step among the token set $\mathcal{V}$, where low entropy means low uncertainty. For example, in code writing, the structure of a code sequence is regularized where few changes can be made [16]. More explicitly, for a typical English pangram such as ‘The quick brown fox jumps over the lazy dog’ [12], both humans and machines should generate similar if not identical output. For example, when provided with the prompt ‘The quick brown fox jumps over the lazy’, the trained LLaMA-2-7B [29] outputs an empirical probability above 0.8 for the next word ‘dog’.
Figure 1: Example of a low-entropy scenario.
Problem modeling. Low-entropy scenarios exist in text generation tasks of LLMs. We aim to model a simple problem protocol for the low-entropy generation scenario. For simplicity, we consider the low-entropy scenario where only one token probability is significantly large. Specifically, denote $p_{max}$ as the largest probability of a token in the probability distribution $P_{M}(\cdot|x^{-N_{p}:(t-1)})$. We make an intuitive assumption that except $p_{max}$, other $|\mathcal{V}|-1$ probabilities are small enough to uniformly fill in the remaining $1-p_{max}$ probability value. Figure 1 demonstrates the low-entropy scenario given the prompt ‘The quick brown fox jumps over the lazy’.
4 Method: Sampling Then Accepting
In this section, we propose the Sampling One Then Accepting (STA-1) method, and discuss detecting the STA-1 generated text using the $z$-test. Theoretically, we show that STA-1 is unbiased and its type II error of the $z$-test has statistical guarantees. Next, we analyze all unbiased watermarks under the low-entropy protocol in Section 3. We finally introduce Sampling M Then Accepting (STA-M), an extension of STA-1.
4.1 Sampling One Then Accepting
We start by proposing the Sampling One Then Accepting (STA-1) method in Algorithm 1, which is always unbiased and easy to analyze. First, the hash value of the last generated token is computed and employed as the seed of a random number generator (RNG). We use the RNG to divide the token set into a green and a red list [12]. Next, we sample from the original LLM output distribution (as depicted in Line 4 of Algorithm 1), accept the sampling if the token is in the green list (as depicted in Line 5 and Line 6 in Algorithm 1), sample again if the token is in the red list (as depicted in Line 7 in Algorithm 1), and the second sampling is always accepted.
Algorithm 1 STA-1 Text Generation
Input: A pretrained LLM $P_{M}$, a key $k\in K$, the proportion of green list $\gamma\in(0,1)$, and a prompt $x^{-N_{p}:0}$
1:  for $t=1,2\dots,T$ do
2:     Get the probability distribution of tokens $p^{t}=P_{M}(\cdot|x^{-N_{p}:(t-1)})$
3:     Compute the hash of the last token $x^{t-1}$. Partition the token set $\mathcal{V}$ to form the green $G$ and red $R$ list based on key $k$, the hash, and the proportion $\gamma$
4:     Sample the candidate token $x_{c}^{t}$ with $p^{t}$
5:     if $x_{c}^{t}\in G$  then
6:        Accept the sampling, the next generated token $x^{t}=x_{c}^{t}$
7:     else
8:        Deny the sampling, sample $x^{t}$ from the distribution $p^{t}$
9:     end if
10:  end for
Output: The generated text $x^{1:T}$
STA-1 is a simple but effective watermarking method. The properties of STA-1 include: (1) STA-1 is an unbiased watermark; (2) The number of green list tokens in STA-1 generated texts has a lower bound on its mean and an upper bound on its variance, which provides statistical guarantees on the detection of STA-1; (3) STA-1 has a lower risk for low-entropy generation compared to previous work. In deriving theoretical results, we assume that the key $k$ is randomly sampled from a uniform distribution. Therefore, the random partition of green and red lists associated with this key is also uniform [12, 11]. We start by analyzing the unbiased characteristic of STA-1.
Theorem 1. The STA-1 method (Algorithm 1) is an unbiased watermark.
Proof. See Appendix B.1.
4.1.1 Statistical test guarantees of STA-1
Detecting the STA-1 generated text. The detection of STA-1 compares the empirical proportion of green list tokens in the given text against the green list proportion $\gamma$ [12]. We employ the $z$-test where the null hypothesis ($H_{0}$) is that the text is generated without knowing the green-red list partition [12]. Denote $|S|_{G}$ as the number of green list tokens in this text. Under $H_{0}$, $|S|_{G}$ follows a Bernoulli distribution $B(T,\gamma)$ with a mean of $\gamma T$ and a variance of $\gamma(1-\gamma)T$. The $z$-score is calculated with the empirical $|S|_{G}$ as
$z=\frac{|S|_{G}-\gamma T}{\sqrt{\gamma(1-\gamma)T}}.$ (3)
The alternative hypothesis ($H_{a}$) is that the text is generated with STA-1. Under $H_{a}$, $|S|_{G}$ is expected to be larger than $\gamma T$. We can detect watermarked texts with a certain confidence level if the $z$-score exceeds a $z$ threshold. For example, if $z>2$, we are more than 97.7% confident that the text is watermarked under the one-tail test.
To ensure the effectiveness of the $z$-test, under $H_{a}$, a lower bound on the expectation of $|S|_{G}$ and an upper bound on the variance of $|S|_{G}$ are required. We establish the necessary lower and upper bounds in the following theorem. Because both bounds are related to the Gini index of the LLM output distribution, we define the Gini index first.
Definition 2 (Gini index). Given a discrete probability distribution $p=(p_{1},p_{2},\cdots,p_{N})$, the Gini index of $p$ is defined as
$Gini(p)=\sum_{i=1}^{N}p_{i}(1-p_{i}).$ (4)
A low Gini index implies less uncertainty in the probability distribution, resulting in a low-entropy scenario. Next, we propose the mean and variance bounds of $|S|_{G}$.
Theorem 2. For STA-1 generated text sequences with $T$ tokens, let the random green list have a fixed size of $\gamma|\mathcal{V}|$, and $p^{t}_{i}$ denote the LLM’s raw output probability of the $i$-th token in $\mathcal{V}$ at step $t$, $i=1,2,\cdots,|\mathcal{V}|$, $p^{t}=(p^{t}_{1},p^{t}_{2},\cdots,p^{t}_{|\mathcal{V}|})$. If an STA-1 generated sequence $S$ has an average Gini index larger than or equal to $Gini^{*}$, that is,
$\frac{1}{T}\sum_{t=1}^{T}Gini(p^{t})=\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{|\mathcal{V}|}p^{t}_{i}(1-p^{t}_{i})\geq Gini^{*}.$
Then the expectation of $|S|_{G}$ is at least
$\mathbb{E}(|S|_{G})\geq\gamma T+(1-\gamma)\gamma TGini^{*}.$ (5)
With one additional assumption that $\gamma$ and $Gini^{*}$ satisfy $\gamma+(1-\gamma)\gamma Gini^{*}\geq 0.5$, the variance of $|S|_{G}$ is at most
$\mathbb{V}(|S|_{G})\leq T[\gamma+(1-\gamma)\gamma Gini^{*}][1-\gamma-(1-\gamma)\gamma Gini^{*}].$ (6)
Proof. See Appendix B.2.
Remark 1. The additional assumption required for the variance upper bound, $\gamma+(1-\gamma)\gamma Gini^{*}\geq 0.5$, implies that a larger green list is necessary in low-entropy scenarios to establish an upper bound on the variance of $|S|_{G}$. By selecting $\gamma\geq 0.5$, this assumption holds for any $Gini^{*}$.
Remark 2. Compared to the Spike entropy proposed by Kirchenbauer et al. [12], the Gini index is a commonly used metric in machine learning to measure the uncertainty of a probability distribution, such as CART decision tree [3].
Example 1. We show an example for a typical $\gamma$. Let $\gamma=0.5$, this bound becomes
$\mathbb{E}(|S|_{G})\geq\frac{1}{2}T+\frac{1}{4}TGini^{*},$ (7)
$\mathbb{V}(|S|_{G})\leq T[\frac{1}{2}+\frac{1}{4}Gini^{*}][\frac{1}{2}-\frac{1}{4}Gini^{*}]=T[\frac{1}{4}-\frac{1}{16}Gini^{*2}].$ (8)
Note that $Gini^{*}$ is the average Gini index. When the generation becomes more uncertain, $Gini^{*}$ increases and we can expect a higher number of green list tokens with a lower variance. Practically in low-entropy scenarios, with probability masses concentrated on one or a few tokens, those tokens are likely to be generated frequently regardless of the green and red list partition in STA-1. Thus, fewer tokens in the green list are expected. This weakens the strength of watermarking methods and makes watermark detection challenging, which is consistent with the theorem.
4.1.2 Discussion on the low-entropy protocol
Previous work claims that unbiased watermarks can avoid the tradeoff between watermark strength and text quality [11]. We challenge this claim by first considering the following example.
Example 2. Assuming that the token set only includes two tokens $\mathcal{V}=\{A,B\}$, at a typical step, an LLM outputs the probability of generating $A$ ($p_{A}$) and $B$ ($p_{B}$) as $(p_{A},p_{B})=(0.8,0.2)$. Consider the following two unbiased watermarks. (1) With a probability of 0.2 always generating $B$ and with a probability of 0.8 always generating $A$. (2) With a probability of 0.5 probabilities becoming $(p_{A},p_{B})=(0.9,0.1)$ and with the other probability of 0.5 becoming $(p_{A},p_{B})=(0.7,0.3)$.
We strongly recommend readers who are unfamiliar with risk-averse or utility theory to Appendix C for a conventional example in finance. In Example 2, one can view the prompt as ‘The quick brown fox jumps over the lazy’, $A$ as the token ‘dog’, and $B$ as all other tokens. It is easy to show that watermarks (1) and (2) are both unbiased. However, risk-averse people [22] will prefer watermark (2) because watermark (2) does not have a probability that only $B$ will be sampled. $B$ represents unsatisfactory outputs in low-entropy scenarios which could significantly harm text quality, and we want the risk of sampling $B$ as low as possible. Again, we refer readers to Appendix C for a better understanding of the analysis. At any generation step, let $x_{max}$ denote the token with the maximum probability $p_{max}$. We measure the risk by the variance [27] of $p_{max}^{w,k}$ among watermark keys, where $p_{max}^{w,k}$ denotes the altered value of $p_{max}$ with a watermarking method and a key $k$. We show that STA-1 has a lower risk compared to previous unbiased watermarks in the following theorem.
Theorem 3. Assume $11-\alpha\leq p_{max}<$ 1, where $\alpha$ represents the partition hyperparameter in Dipmark. For the low-entropy protocol in Section 3, the STA-1 method has a lower variance in the probability of generating $x_{max}$ compared to other unbiased methods (including Dipmark and $\gamma$-reweight) [11, 33]. Formally,
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]<\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right],$ (9)
for any $\alpha\in[0,0.5]$ used in Dipmark, where $p_{max}^{w,k}$ denotes the adjusted probability of the token $x_{max}$ under the respective watermarking method with a key $k\in K$.
Proof. See Appendix B.3.
Example 3. We show a numerical example with $p_{max}=0.8$. For STA-1, based on the proof of the theorem, if the proportion of the green list is 0.5, $\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]=6.4\times 10^{-3}$. For Dipmark ($\alpha\in[0.2,0.5]$) and $\gamma$-reweight, the variance is $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]=\frac{1}{75}\approx 0.013$.
4.2 Sampling M Then Accepting
A low-entropy scenario indicates a low Gini index which weakens the watermark strength based on Theorem 2. To enhance the watermark strength, we propose the Sampling M Then Accepting (STA-M) method, an extension of STA-1. STA-M employs a heuristic threshold $\tau$ for entropy at each generation step. In detail, at generation step $t$, we first calculate the entropy $\tau^{t}$ of the probability distribution $P_{M}(\cdot|x^{-N_{p}:(t-1)})$. If it shows low entropy $\tau^{t}\leq\tau$, we apply STA-1 at this generation step; if it shows high entropy $\tau^{t}>\tau$, we repeat sampling if the previously sampled token is in the red list, and the procedure repeats at most $M$ times. The detailed algorithm and analysis of STA-M can be found in Appendix D.
5 Experiments
In this section, we conducted computational experiments to evaluate the performance of STA-1 and STA-M using two public datasets. We benchmarked our methods against various watermarking baselines on text quality and watermark strength. Moreover, we discussed the variance of generation in the low-entropy dataset. Finally, we conducted a robustness analysis of STA against different watermarking attacks.
5.1 Experimental Setup
Datasets and metrics. We employed two public datasets: C4 subset [24, 12] for news-like text generation and HumanEval [4] for code generation. We evaluated the performance of different watermarking methods on text quality and watermark strength. For watermark strength, we set the $z$ threshold as 2 and 2.5 and report the F1-score and AUC of watermark detection. For text quality, we measured perplexity (PPL) and coherence [9] for generations on C4; We computed PPL and pass@$k$ scores of code generations [4] for HumanEval. We refer readers to Appendix E.1 for more dataset details and the prompt used in datasets.
Baselines. We chose KGW as the biased watermark baseline [12], $\gamma$-reweight [11] and Dipmark [33] as the unbiased watermark baselines. Specifically, we set KGW with a fixed green list proportion $\gamma=0.5$ and diverse logit increments $\delta\in\{1,1.5,2\}$. The partition parameter of Dipmark was set as $\alpha\in\{0.3,0.4,0.5\}$. When $\alpha=0.5$, we report this result as $\gamma$-reweight. Note that $\gamma$-reweight [11] does not include a $z$-test. We implemented the $z$-score in Dipmark [33] by counting the number of tokens in the latter portion of the token set. We also show results without watermarking techniques.
Implementation details. We utilized different variants of LLaMA-2-7B [29] as our generative models, and LLaMA-2-13B to compute perplexity. For hyperparameters in STA-M, we set $M\in\{4,8,16\}$ and two entropy thresholds $\tau$ for different datasets. We conducted a robustness check on $\tau$ in Appendix E.2 and selected different $\tau$s for different datasets in the final experiment. We refer readers to Appendix E.1 for more details on implementation.
Table 1: Result Comparison between Our Methods and Baselines on Text Quality and Watermark Strength for the C4 Dataset.
Text Quality Watermark Strength
$z=2.0$ $z=2.5$
Method PPL Coherence F1 AUC F1 AUC
No Watermark 7.409 0.605 0.049 0.495 0.012 0.499
KGW($\delta$=1) 7.617 0.602 0.956 0.957 0.928 0.932
KGW($\delta$=1.5) 7.870 0.600 0.982 0.982 0.990 0.990
KGW($\delta$=2) 8.221 0.602 0.980 0.980 0.993 0.993
Dipmark($\alpha$=0.3) 7.451 0.604 0.923 0.924 0.893 0.901
Dipmark($\alpha$=0.4) 7.451 0.604 0.951 0.951 0.943 0.945
$\gamma$-reweight 7.437 0.611 0.969 0.969 0.964 0.965
STA-1 7.343 0.600 0.957 0.956 0.961 0.961
STA-4($\tau$=1.35) 7.735 0.588 0.981 0.981 0.991 0.991
STA-8($\tau$=1.35) 8.161 0.589 0.969 0.968 0.991 0.991
STA-16($\tau$=1.35) 7.988 0.585 0.974 0.974 0.992 0.992
5.2 Results on C4
For the C4 dataset, each method generates at least 500 text sequences with at least $200\pm 5$ tokens [12]. Table 1 demonstrates the text quality and watermark strength for each method, and we present generated text examples in Appendix E.3. As depicted in Table 1, the proposed STA-1 method achieves comparable perplexity and coherence compared to no watermark generation. The text quality results are consistent with other unbiased watermarks including Dipmark and $\gamma$-reweight, showing STA-1 is also unbiased empirically. In terms of watermark strength, STA-M ($M\in\{4,8,16\}$) outperforms unbiased watermarks including STA-1 and Dipmark, and has similar watermark strength as biased watermarks KGW ($\delta\in\{1.5,2\}$). Overall in the high-entropy generation task, the unbiased STA-1 method is comparable to other unbiased watermarks; The STA-M method can improve the watermark strength by sacrificing minor text quality.
Table 2: Result Comparison between Our Methods and Baselines on Text Quality and Watermark Strength for the HumanEval Dataset.
Text Quality Watermark Strength
$z=2.0$ $z=2.5$
Method PPL Pass@$1$ Pass@$5$ Pass@$10$ F1 AUC F1 AUC
No Watermark 3.275 0.147 0.405 0.537 0.114 0.494 0.072 0.497
KGW($\delta$=1) 3.288 0.119 0.326 0.415 0.471 0.643 0.416 0.627
KGW($\delta$=1.5) 3.285 0.104 0.308 0.427 0.720 0.770 0.650 0.730
KGW($\delta$=2) 3.650 0.077 0.254 0.372 0.757 0.795 0.733 0.785
Dipmark($\alpha$=0.3) 3.081 0.142 0.392 0.512 0.518 0.665 0.423 0.625
Dipmark($\alpha$=0.4) 3.020 0.135 0.393 0.512 0.516 0.668 0.429 0.634
$\gamma$-reweight 3.314 0.130 0.371 0.488 0.522 0.671 0.479 0.655
STA-1 3.006 0.155 0.394 0.494 0.526 0.633 0.442 0.611
STA-4($\tau$=1.95) 3.175 0.151 0.392 0.500 0.633 0.685 0.594 0.679
STA-8($\tau$=1.95) 2.842 0.148 0.399 0.537 0.652 0.703 0.587 0.675
STA-16($\tau$=1.95) 3.024 0.145 0.382 0.476 0.725 0.764 0.640 0.717
5.3 Results on HumanEval
In this section, we compare our methods against baselines on the HumanEval dataset. Table 2 presents the perplexity, pass@$k$ scores, and watermark strength for all methods. First, we focus on the result analysis for all unbiased watermarks. As reported, our STA-1 method achieves similar perplexity and pass@$k$ scores compared to no watermarking and other unbiased watermarking methods.
Table 3: Comparison on the Risk of Unsatisfactory Outputs for Unbiased Watermarks. For space concern, we denote the number of passed problems as PP, the number of passed codes as PC, and the average number of passed codes per passed problem as PC per PP (PC/PP).
Method PPL Variance PP PC PC per PP
Dipmark($\alpha$=0.3) 1.535 84 233 2.675
Dipmark($\alpha$=0.4) 1.853 84 221 2.631
$\gamma$-reweight 1.722 80 214 2.774
STA-1 1.461 81 254 3.136
Moreover, we examine the risk of unsatisfactory outputs produced by unbiased watermarks for low-entropy generations. Specifically, we compare different unbiased watermarks in terms of four more metrics. We ran 10 times of code generation for each problem using different unbiased watermarking methods with 10 different keys. Table 3 reports the variance of perplexity, the number of passed problems, [^2] If the problem is solved by any one generation out of 10 runs, it is considered passed. the number of passed codes, and the average number of passed codes among all passed problems. In particular, the STA-1 method demonstrates the lowest variance of perplexity compared to Dipmark($\alpha$=0.3), Dipmark($\alpha$=0.4), and $\gamma$-reweight with a variance of 1.461 compared to 1.535, 1.853, and 1.722, respectively. A lower variance indicates a lower risk among different text generations under different keys. Additionally, we show the average number of passed codes among all passed problems. For example, 3.136 in Table 3 means among all solved problems, an average of 3.136 generated codes are accurate w.r.t. 10 generations by STA-1. We conclude from Table 3 that although Dipmark solves more problems, it fails to provide consistent accurate codes among different generations. Instead, our method outperforms other unbiased watermarks (Dipmark($\alpha$=0.3), Dipmark($\alpha$=0.4), $\gamma$-reweight) in providing consistency, with an average number of passed codes of 3.136 compared to 2.675, 2.631, and 2.774, respectively. In summary, the STA-1 method has a lower risk when generating low-entropy texts, as discussed in Theorem 3.
In terms of watermark strength, STA-M ($M\in\{4,8,16\}$) yields higher watermark strength in comparison to all unbiased watermarks while maintaining similar pass scores. The STA-16 method achieves comparable watermark strength against biased watermark KGW($\delta=2$) with an AUC of 0.764 ($z=2$) against 0.795. The text quality is maintained with a pass@$10$ of 0.476, highlighting the efficacy of the heuristics to enhance watermark strength at high-entropy generation steps.
Table 4: Attacking Watermarks for the C4 Dataset.
        Attack Setting No Attack GPT-3.5 DIPPER-1 DIPPER-2 Copy-Paste
Method F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC
KGW($\delta=1$) 0.96 0.96 0.27 0.57 0.13 0.53 0.15 0.54 0.68 0.75
KGW($\delta=1.5$) 0.98 0.98 0.41 0.62 0.22 0.56 0.27 0.57 0.90 0.90
KGW($\delta=2$) 0.98 0.98 0.54 0.68 0.30 0.58 0.40 0.62 0.95 0.95
Dipmark($\alpha=0.3$) 0.92 0.92 0.29 0.57 0.24 0.55 0.26 0.55 0.61 0.70
Dipmark($\alpha=0.4$) 0.95 0.95 0.38 0.61 0.31 0.58 0.34 0.59 0.75 0.79
$\gamma$-reweight 0.97 0.97 0.41 0.61 0.32 0.57 0.36 0.60 0.74 0.78
STA-1 0.96 0.96 0.47 0.63 0.39 0.60 0.46 0.63 0.78 0.81
STA-4($\tau$=1.35) 0.98 0.98 0.72 0.78 0.65 0.73 0.69 0.75 0.95 0.95
STA-8($\tau$=1.35) 0.97 0.97 0.78 0.81 0.71 0.77 0.76 0.79 0.95 0.95
STA-16($\tau$=1.35) 0.97 0.97 0.76 0.80 0.68 0.74 0.78 0.81 0.95 0.95
5.4 Attacking STA
We assessed the robustness of different watermarking methods under different attacks consisting of paraphrasing using GPT-3.5, two configurations of the DIPPER attack [14], and the copy-paste attack [12]. Detailed settings of different attacks are described in Appendix E.4. Table 4 reports the F1-score and AUC of watermark detection under each attack with $z=2$. As reported, STA-M is robust against different attacks. The reason is that GPT-3.5 and DIPPER are both LLM-based attacks that sample tokens to replace tokens in given texts. However, LLM-based attacks can hardly replace tokens in STA-M generated text because of the high text quality, and cannot find enough tokens to replace them because of the high watermark strength of STA-M. Meanwhile, the high watermark strength ensures robustness against the copy-paste attack.
6 Conclusions and Future Work
In this work, we propose a new unbiased watermarking method named STA-1. We clarify the text quality (regarding the risk of unsatisfactory outputs) and watermark strength tradeoff of unbiased watermarks in low-entropy scenarios. We also extend STA-1 to STA-M which can enhance watermark strength with little text quality shifts. Experimental results on low-entropy datasets prove that STA-1 is comparable to other unbiased watermarks and has a low risk. Moreover, the results from both datasets demonstrate the efficacy and robustness of STA-M. Future work of our study can be categorized into several ways. First, watermarking low-entropy tasks is still challenging and future work can devise better watermarking methods. Second, future work could incorporate more datasets and generative LLMs for evaluation of our method.
7 Related Work
With the development of LLMs, the idea of watermarking LLMs has been proposed [1, 12] and widely explored [18]. Existing white-box watermarking techniques can be categorized into watermarking during logits and probabilities generation [13, 16, 11, 30, 8, 35, 34, 25, 28], and watermarking by sampling candidate tokens [5, 15, 10]. We refer readers to Appendix F for a detailed discussion on related work.
Appendix A Details of Previous Methods
Distribution reweighting refers to methods that adjust the output distribution $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ at each step $t$ by artificially increasing probabilities for certain tokens while reducing those for others. The direction and magnitude (increasing or decreasing) of change in probability mass for a token are determined by the private key $k$.
KGW [12] first randomly splits the vocabulary set $\mathcal{V}$ into two non-overlapping lists based on a uniformly distributed key $k$: a ‘green’ list and a ‘red’ list. This method has two versions: the ‘hard’ version completely ignores the red list tokens and only samples tokens from the green list; The ‘soft’ version adds a predefined constant $\delta$ to logits of green list tokens while keeping logits of red list tokens fixed. The soft KGW reweights distribution as
$P_{M,w}(x^{t}=j|x^{-N_{p}:(t-1)};k)=\frac{\exp\left(l_{j}^{t}+\mathds{1}_{\text{Green}}(j)\delta\right)}{\sum_{i\in\text{Red}}\exp(l_{i}^{t})+\sum_{i\in\text{Green}}\exp(l_{i}^{t}+\delta)},$
where $j$ denotes the $j$-th token within the vocabulary set, $l_{j}^{t}$ is its logit output by the original LLM at step $t$, and $\mathds{1}_{\text{Green}}(j)$ is an indicator function having a value of 1 when $j$ is in the green list and 0 otherwise.
Wu et al. [33] propose an unbiased reweighting method, named Dipmark. Dipmark arranges all probability masses over the vocabulary set from the original LLM output consecutively within the interval $[0,1]$ and then randomly permutes their orders based on a key $k$. A hyperparameter $\alpha\in[0,0.5]$ partitions the probability interval $[0,1]$ into three segments: $[0,\alpha]$, $(\alpha,1-\alpha]$, and $(1-\alpha,1]$. Probability masses in the first segment are set to 0, those in the second remain constant, and those in the third are doubled. Denote the token order after permutation as $\widetilde{\mathcal{V}}$, the adjusted probability for the $j$-th token within $\widetilde{\mathcal{V}}$ is $P_{M,w}(x^{t}=j|x^{-N_{p}:(t-1)};k)=F(j|\widetilde{\mathcal{V}})-F(j-1|\widetilde{\mathcal{V}})$, with $F(j|\widetilde{\mathcal{V}})$ being defined as
$F(j|\widetilde{\mathcal{V}})=\max\left[\sum_{i\in\widetilde{\mathcal{V}}:i\leq j}P_{M}(x^{t}=i|\cdot)-\alpha,0\right]+\max\left[\sum_{i\in\widetilde{\mathcal{V}}:i\leq j}P_{M}(x^{t}=i|\cdot)-(1-\alpha),0\right].$
Appendix B Proofs
B.1 Proof of Theorem 1
To simplify notation, we denote the size of the vocabulary set $|\mathcal{V}|$ as $N$, the size of the green list as $N_{G}$, and the size of the red list as $N_{R}$. Given the proportion of green list $\gamma$, we have $N_{G}=\gamma N$ and $N_{R}=(1-\gamma)N$. At a generation step, let $p=(p_{1},p_{2},\cdots,p_{N})$ denote the raw probability output by the LLM over the vocabulary set. Let $j$ represent a token within the vocabulary set, $j\in(1,2,\cdots,N)$. We denote by $p_{j}^{w,k}$ the adjusted probability of token $j$ under the STA-1 watermarking method with key $k$. The key $k$ is sampled randomly from a uniform distribution $P_{K}(k)$.
To conveniently compute $\mathbb{E}_{k\sim P_{K}(k)}\left[p_{j}^{w,k}\right]$, we consider the uniformly random partition of green and red lists associated with the uniformly distributed key $k$ as the following process. Initially, token $j$ is randomly assigned to the green list with a probability of $\gamma$ and to the red list with a probability of $1-\gamma$. Subsequently, tokens are randomly sampled from the remaining pool to fill the green list, with all remaining tokens then placed in the red list. For the adjusted probability, we have
$p_{j}^{w,k}=\begin{cases}p_{j}+\left(\sum_{i\in R}p_{i}\right)p_{j}&j\in G\\ \left(\sum_{i\in R}p_{i}\right)p_{j}&j\in R\end{cases}.$
Next, we first analyze the scenario where $j\in G$ and compute $\mathbb{E}_{G,R:j\in G}\left[p_{j}^{w,k}\right]$. The expectation is taken over uniformly random partitions of green/red lists that fulfill $j\in G$. Let
$h_{j}(p)=\mathbb{E}_{G,R:j\in G}\left[p_{j}^{w,k}\right]=\mathbb{E}_{G,R:j\in G}\left[p_{j}+\left(\sum_{i\in R}p_{i}\right)p_{j}\right].$
Note that $h_{j}(p)$’s value remains unchanged under permutations in the order of the remaining tokens $\{p_{i},i\neq j\}$. Thus, we have the equality that $h_{j}(p)=\mathbb{E}_{\Pi}\left[h_{j}(\Pi p_{-j})\right]$, where $\Pi$ represents a random permutation of the remaining tokens $p_{-j}$ while preserving the position of $p_{j}$. Since $h_{j}(\Pi p_{-j})$ is a linear function of $p_{-j}$, we then get
$h_{j}(p)=\mathbb{E}_{\Pi}\left[h_{j}(\Pi p_{-j})\right]=h_{j}\left(\mathbb{E}_{\Pi}\left[\Pi p_{-j}\right]\right).$
The expectation of the probability values at the remaining $(N-1)$ positions over permutations of their corresponding tokens $\mathbb{E}_{\Pi}\left[\Pi p_{-j}\right]$ yields a probability distribution $\bar{p}$ where $\bar{p}_{j}=p_{j}$ and $\bar{p}_{i}=(1-p_{j})/(N-1)$ for $i\neq j$. With this $\bar{p}$, we derive that
$h_{j}(p)=h_{j}(\bar{p})$ $=\mathbb{E}_{G,R:j\in G}\left[\bar{p}_{j}+\left(\sum_{i\in R}\bar{p}_{i}\right)\bar{p}_{j}\right]$
$=p_{j}+\frac{N_{R}}{N-1}(1-p_{j})p_{j}.$
Then, we analyze the scenario where $j\in R$ and compute $\mathbb{E}_{G,R:j\in R}\left[p_{j}^{w,k}\right]$. Let
$f_{j}(p)=\mathbb{E}_{G,R:j\in R}\left[p_{j}^{w,k}\right]=\mathbb{E}_{G,R:j\in R}\left[\left(\sum_{i\in R}p_{i}\right)p_{j}\right].$
For the same reasons as illustrated above and using the same definition of $\bar{p}$, we have
$f_{j}(p)=f_{j}(\bar{p})$ $=\mathbb{E}_{G,R:j\in R}\left[\left(\sum_{i\in R}\bar{p}_{i}\right)\bar{p}_{j}\right]$
$=\left(p_{j}+\frac{(N_{R}-1)(1-p_{j})}{(N-1)}\right)p_{j}$
$=p_{j}^{2}+\frac{(N_{R}-1)}{N-1}(1-p_{j})p_{j}.$
Finally, combining the random partition process of green and red lists described at the beginning of the proof with the derived expressions for $h_{j}(p)$ and $f_{j}(p)$, we obtain that
$\mathbb{E}_{k\sim P_{K}(k)}\left[p_{j}^{w,k}\right]$ $=\gamma h_{j}(p)+(1-\gamma)f_{j}(p)$
$=\gamma p_{j}+\gamma\frac{N_{R}}{N-1}(1-p_{j})p_{j}+(1-\gamma)p_{j}^{2}+(1-\gamma)\frac{(N_{R}-1)}{N-1}(1-p_{j})p_{j}$
$=\left(\gamma+\frac{N_{R}-(1-\gamma)}{N-1}\right)p_{j}+\left((1-\gamma)-\frac{N_{R}-(1-\gamma)}{N-1}\right)p_{j}^{2}$
$=p_{j},$
with $N_{R}=(1-\gamma)N$. This concludes the proof.
B.2 Proof of Theorem 2
In this proof, we employ the notations introduced in the proof of Theorem 1 in Section B.1, and we leverage the results derived from that theorem’s proof.
For a token $j$ within the vocabulary set, $j\in(1,2,\cdots,N)$, we consider the identical random partition process of green and red lists as described at the beginning of the proof of Theorem 1. If $j$ is initially assigned to the green list, according to the proof of Theorem 1, its expected adjusted probability over uniformly random green/red list partitions that fulfill $j\in G$ satisfies
$\mathbb{E}_{G,R:j\in G}\left[p_{j}^{w,k}\right]$ $=p_{j}+\frac{N_{R}}{N-1}(1-p_{j})p_{j}$
$=p_{j}\left[\frac{N-1+N_{R}(1-p_{j})}{N-1}\right]$
$\geq p_{j}\left[\frac{N+N_{R}(1-p_{j})}{N}\right]$
$=p_{j}+(1-\gamma)p_{j}(1-p_{j}),$
where the inequality holds because the denominator is less than the numerator, and adding 1 to both leads to a decrease in the value.
Recall that each token within the vocabulary set has a probability of $\gamma$ being assigned to the green list. Thus, the overall probability of sampling a token from the green list has the lower bound
$\mathbb{P}(G)\coloneqq\mathbb{P}(\text{sampling a token}\in G)$ $=\sum_{j=1}^{N}\gamma\mathbb{E}_{G,R:j\in G}\left[p_{j}^{w,k}\right]$
$\geq\gamma\sum_{j=1}^{N}p_{j}+(1-\gamma)p_{j}(1-p_{j})$
$=\gamma+\gamma(1-\gamma)\sum_{j=1}^{N}p_{j}(1-p_{j}).$
Note that this lower bound applies to every generation step $t$. Let $p^{t}$ denote the LLM’s original output probability distribution at step $t$, and $G^{t}$ denote the event of sampling a token from the green list at step $t$, we then have
$\mathbb{P}(G^{t})\geq\gamma+\gamma(1-\gamma)\sum_{j=1}^{N}p_{j}^{t}(1-p_{j}^{t})=\gamma+\gamma(1-\gamma)Gini(p^{t}).$
It is important to highlight that this lower bound holds significant meaning, as it strictly exceeds the naive lower bound for $\mathbb{P}(G^{t})$, which is $\gamma$. This bound serves as a crucial element in the proof of Theorem 2. For the expectation of the number of green list tokens in the sequence, we can derive that
$\mathbb{E}(|S|_{G})$ $=T\mathbb{E}_{t}\left[\mathbb{P}(G^{t})\right]\geq T\mathbb{E}_{t}\left[\gamma+\gamma(1-\gamma)Gini(p^{t})\right]$
$\geq T\left[\gamma+\gamma(1-\gamma)Gini^{*}\right]=\gamma T+(1-\gamma)\gamma TGini^{*},$
where the lower bound $Gini^{*}$ for the average Gini index is provided as a condition in the theorem.
Next, regarding the variance of $|S|_{G}$, it is worth noting that the success of sampling a token from the green list at each step $t$ can be viewed as a Bernoulli random variable with a success probability of $\mathbb{P}(G^{t})$. This Bernoulli random variable has a variance of $\mathbb{P}(G^{t})[1-\mathbb{P}(G^{t})]$. The sum of these Bernoulli random variables across all $T$ steps gives us $|S|_{G}$. Because these random variables are independent of each other, the variance of their sum equals the sum of their variances. Consequently, we can obtain that
$\mathbb{V}(|S|_{G})$ $=T\mathbb{E}_{t}\left[\mathbb{P}(G^{t})[1-\mathbb{P}(G^{t})]\right]$
$\leq T\mathbb{E}_{t}[\mathbb{P}(G^{t})]\left[1-\mathbb{E}_{t}[\mathbb{P}(G^{t})]\right]$
$\leq T\left[\gamma+(1-\gamma)\gamma Gini^{*}\right]\left[1-\gamma-(1-\gamma)\gamma Gini^{*}\right],$
where the first inequality holds by applying Jensen’s inequality to a concave function of $\mathbb{P}(G^{t})$, and the second inequality is valid because 1) $\mathbb{E}_{t}\left[\mathbb{P}(G^{t})\right]\geq\gamma+(1-\gamma)\gamma Gini^{*}$ as shown above; 2) the function $x(1-x)$ is decreasing in the range $x\in[0.5,1]$; and 3) it is assumed in the theorem that $\gamma+(1-\gamma)\gamma Gini^{*}\geq 0.5$. This concludes the proof.
B.3 Proof of Theorem 3
In this proof, we continue utilizing the notations introduced in the proof of Theorem 1 in Section B.1.
We start with the variance calculation for the STA-1 method. Because STA-1 is an unbiased watermark by Theorem 1, we have $\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]=\mathbb{E}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$. Considering the identical uniformly random partition process of green and red lists associated with the uniformly distributed key $k$ as in the proof of Theorem 1, depending on whether the token $x_{max}$ is assigned to the green list or not initially, $p_{max}^{w,k}$ have two possible realizations:
$p_{max}^{w,k}=\begin{cases}p_{max}+\left(\sum_{i\in R}p_{i}\right)p_{max}&x_{max}\in G\\ \left(\sum_{i\in R}p_{i}\right)p_{max}&x_{max}\in R\end{cases}.$
Under the assumption that the probabilities of the other $N-1$ tokens uniformly fill in the remaining $(1-p_{max})$ probability mass, each $p_{i}$, $i\in(1,2,\cdots,N)\text{ and }i\neq x_{max}$, equals $(1-p_{max})/(N-1)$. Therefore, if $x_{max}\in G$, $p_{max}^{w,k}=p_{max}+N_{R}(1-p_{max})p_{max}/(N-1)$, and this value is fixed for all partitions of green/red lists that fulfill $x_{max}\in G$. Then we have
$\mathbb{E}_{G,R:x_{max}\in G}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]=\left[\frac{N_{R}(1-p_{max})p_{max}}{(N-1)}\right]^{2}.$
Similarly, if $x_{max}\in R$, we get
$\mathbb{E}_{G,R:x_{max}\in R}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]=\left[\left(\frac{(N_{R}-1)(1-p_{max})}{N-1}+p_{max}\right)p_{max}-p_{max}\right]^{2}.$
With these two expected values, and recalling that $x_{max}$ has a probability of $\gamma$ of being assigned to the green list and a probability of $1-\gamma$ of being assigned to the red list, the variance for the STA-1 method is
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]=\mathbb{E}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$
$=\gamma\mathbb{E}_{G,R:x_{max}\in G}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]+(1-\gamma)\mathbb{E}_{G,R:x_{max}\in R}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$
$=\gamma\left[\frac{N_{R}(1-p_{max})p_{max}}{(N-1)}\right]^{2}+(1-\gamma)\left[\left(\frac{(N_{R}-1)(1-p_{max})}{N-1}+p_{max}\right)p_{max}-p_{max}\right]^{2}$
$=p_{max}^{2}(1-p_{max})^{2}\left[\gamma\frac{N_{R}^{2}}{(N-1)^{2}}+(1-\gamma)\frac{N_{G}^{2}}{(N-1)^{2}}\right]$
$=p_{max}^{2}(1-p_{max})^{2}\gamma(1-\gamma)\frac{N^{2}}{(N-1)^{2}}.$
Next, we compute the variance for the Dipmark method with a partition hyperparameter $\alpha$. Note that $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{E}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$ holds because Dipmark is also unbiased. In Dipmark, the uniformly distributed key $k$ controls the randomness of permutations. Under the same assumption that $p_{i}=(1-p_{max})/(N-1)$ for $i\neq x_{max}$, the relative orders among these $(N-1)$ tokens become irrelevant in the permutation. Therefore, there are a total of $N$ unique permutations, each with a probability of $1/N$. Specifically, in the first unique permutation, there are 0 tokens $i$ where $i\neq x_{max}$ placed to the left of $x_{max}$ and $(N-1)$ tokens $i$ where $i\neq x_{max}$ placed to the right of $x_{max}$. In the second one, there is 1 token on the left and $(N-2)$ tokens on the right, and so forth. The last permutation has $(N-1)$ tokens on the left and 0 on the right. If $j$ such tokens are on the left of $x_{max}$, $j=0,1,\cdots,(N-1)$, the corresponding $p_{max}^{w,k}$ is
$p_{max}^{w,k}=2p_{max}-1+2j\frac{(1-p_{max})}{(N-1)},$
given that $11-\alpha\leq p_{max}<$ 1 as assumed in the condition. Therefore, the variance for the Dipmark method with a partition hyperparameter $\alpha$ is
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]$ $=\mathbb{E}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$
$=\frac{1}{N}\sum_{j=0}^{N-1}\left[p_{max}-1+2j\frac{(1-p_{max})}{(N-1)}\right]^{2}$
$=-(p_{max}-1)^{2}+\frac{1}{N}\sum_{j=0}^{N-1}4j^{2}\frac{(1-p_{max})^{2}}{(N-1)^{2}}$
$=(1-p_{max})^{2}\frac{(N+1)}{3(N-1)}.$
Note that, this variance value does not depend on $\alpha$. When $\alpha=0.5$, Dipmark becomes $\gamma$-reweight. Therefore, $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]$.
To compare $\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]$ and $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]$, consider
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]$ $=p_{max}^{2}(1-p_{max})^{2}\gamma(1-\gamma)\frac{N^{2}}{(N-1)^{2}}$
$<\frac{1}{4}(1-p_{max})^{2}\frac{N^{2}}{(N-1)^{2}}$
$=(1-p_{max})^{2}\frac{(N+1)}{3(N-1)}\times\frac{3}{4}\frac{N^{2}}{N^{2}-1},$
where $N^{2}/(N^{2}-1)$ is a decreasing function on $N$ and $N^{2}/(N^{2}-1)<4/3$ for $N>2$. Therefore, for a real-world vocabulary set where $N\gg 2$, we have
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]<(1-p_{max})^{2}\frac{(N+1)}{3(N-1)}=\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right],$
which concludes the proof.
Appendix C Example of Risk-averse
St. Petersburg paradox [31]. Assume that one must choose either one lottery from the following two lotteries. (1) Lottery 1 ($L1$) has a 0.8 probability of earning nothing and the other 0.2 probability of losing 1,000 dollars. (2) Lottery 2 ($L2$) has a 0.5 probability of losing 100 dollars and the other 0.5 probability of losing 300 dollars.
It is easy to show that $L1$ and $L2$ have the same expected outcome that $0.8\times 0-0.2\times 1000=-0.5\times 100-0.5\times 300=-200$. However, risk-averse people will choose $L2$ as they do not want to take the risk of losing 1,000 dollars.
Computationally, assume the person has 1,001 dollars in total and the utility function is $\ln(Y)$ [6], where $Y$ is the wealth. The utility function measures happiness. It is a concave function (such as $\ln(Y)$) because people are happier if they are wealthier ($\ln^{\prime}(Y)>0$) but the increment of happiness decreases as the wealth increases ($\ln^{\prime\prime}(Y)<0$).
The weighted utility of $L1$ and $L2$ are as follows
$U(L1)=0.8\times\ln(1001)+0.2\times\ln(1)\approx 5.53,$
$U(L2)=0.5\times\ln(901)+0.5\times\ln(701)\approx 6.68.$
Based on the weighted utility, risk-averse people will choose $L2$.
Link the lottery example to Example 2 in Section 4.1.2. Because of the low-entropy setting, sampling $B$ results in a huge loss in text quality. Suppose we treat sampling $A$ as earning nothing and sampling $B$ as losing 1,000 for text quality. In this case, we should minimize the risk of sampling $B$. Also in this case, the two unbiased watermarks in Example 2 can be viewed as $L1$ and $L2$ in the lottery example. Sampling $B$ may not be a big issue in high-entropy scenarios because it should not significantly harm text quality as much as 1,000.
Algorithm 2 STA-M Text Generation
Input: A pretrained LLM $P_{M}$, a key $k\in K$, the proportion of green list $\gamma\in(0,1)$, the number of maximum samples per step $M$, a entropy threshold $\tau$, and a prompt $x^{-N_{p}:0}$
1:  for $t=1,2,\dots,T$ do
2:     Get the probability distribution of tokens $p^{t}=P_{M}(\cdot|x^{-N_{p}:(t-1)})$
3:     Compute the entropy $\tau^{t}$ of $p^{t}$
4:     if $\tau^{t}<\tau$ then
5:        $M^{t}=1$
6:     else
7:        $M^{t}=M$
8:     end if
9:     Compute the hash of the last token $x^{t-1}$. Partition the token set $\mathcal{V}$ to form the green $G$ and red $R$ list based on key $k$, the hash, and the proportion $\gamma$
10:     Initialize sample number $m=1$
11:     while $m\leq M^{t}$ and the next token $x^{t}$ not defined  do
12:        Sample the candidate token $x_{c,m}^{t}$ with $p^{t}$
13:        if $x_{c,m}^{t}\in G$ then
14:           Accept the sampling, the next generated token $x^{t}=x_{c,m}^{t}$
15:        else
16:           $m\leftarrow m+1$
17:        end if
18:     end while
19:     if the next token $x^{t}$ not defined then
20:        Sample $x^{t}$ from the distribution $p^{t}$
21:     end if
22:  end for
Output: The generated text $x^{1:T}$
Appendix D STA-M Details
The detailed algorithm of STA-M is shown in Algorithm 2.
Remark 3. STA-M is not unbiased.
We provide a counterexample to show that STA-M is biased. Assume that the vocabulary set consists of four tokens $\{a,b,c,d\}$, and at a generation step, the raw probabilities output by the LLM for these tokens are $\{p_{a}=1/2,p_{b}=1/3,p_{c}=p_{d}=1/12\}$. The proportion of green list $\gamma$ equals $0.5$. Therefore, with a key $k$, two tokens are randomly assigned to the green list, and the red list contains the other two. For the uniformly distributed key $k$, there are six possible random partitions of green and red lists: $\{a,b\in G;c,d\in R\}$, $\{a,c\in G;b,d\in R\}$, $\{a,d\in G;b,c\in R\}$, $\{b,c\in G;a,d\in R\}$, $\{b,d\in G;a,c\in R\}$, and $\{c,d\in G;a,b\in R\}$, each with a probability of $1/6$. Next, considering the token $a$, its adjusted probability under the STA-M watermarking method for each of the six partitions is:
$p_{a}^{w,k}=\begin{cases}\frac{1}{2}+\frac{1}{6}\times\frac{1}{2}+(\frac{1}{6})^{2}\times\frac{1}{2}+\cdots+(\frac{1}{6})^{M}\times\frac{1}{2}&\{a,b\in G;c,d\in R\}\\ \frac{1}{2}+\frac{5}{12}\times\frac{1}{2}+(\frac{5}{12})^{2}\times\frac{1}{2}+\cdots+(\frac{5}{12})^{M}\times\frac{1}{2}&\{a,c\in G;b,d\in R\}\\ \frac{1}{2}+\frac{5}{12}\times\frac{1}{2}+(\frac{5}{12})^{2}\times\frac{1}{2}+\cdots+(\frac{5}{12})^{M}\times\frac{1}{2}&\{a,d\in G;b,c\in R\}\\ (\frac{7}{12})^{M}\times\frac{1}{2}&\{b,c\in G;a,d\in R\}\\ (\frac{7}{12})^{M}\times\frac{1}{2}&\{b,d\in G;a,c\in R\}\\ (\frac{5}{6})^{M}\times\frac{1}{2}&\{c,d\in G;a,b\in R\}\\ \end{cases}.$
With these adjusted probability values, the expectation of the adjusted probability over the six possible partitions is easily derived as
$\mathbb{E}_{k\sim P_{K}(k)}\left[p_{a}^{w,k}\right]$ $=\frac{1}{12}\left[\frac{6}{5}\left(1-(\frac{1}{6})^{M+1}\right)+2\times\frac{12}{7}\left(1-(\frac{5}{12})^{M+1}\right)+2\times(\frac{7}{12})^{M}+(\frac{5}{6})^{M}\right]$
$=\frac{27}{70}-\frac{1}{10}(\frac{1}{6})^{M+1}-\frac{2}{7}(\frac{5}{12})^{M+1}+\frac{1}{6}(\frac{7}{12})^{M}+\frac{1}{12}(\frac{5}{6})^{M},$
which equals $p_{a}=1/2$ only when $M=1$ and is less than $1/2$ for $M\geq 2$. Hence, this counterexample demonstrates that the STA-M method is biased.
Appendix E Experiment
E.1 Experimental Setup
Datasets and metrics. We employed two public datasets which are C4 subset [24, 12] for news-like text generation and HumanEval [4] for code generation. Specifically, C4 represents the high-entropy generation task and HumanEval represents the low-entropy generation task.
C4: We extracted random text segments from the news-like subset of the C4 dataset [24] following Kirchenbauer et al. [12]. For each segment, we removed a fixed number of tokens from the end and the removed tokens served as a ‘baseline’ completion. The remaining tokens were used as the prompt.
HumanEval: HumanEval includes 164 Python problems with test cases and solutions written by humans. We prompted the LLM with these problems. In particular, the prompt was devised as ‘Below is an instruction that describes a task. Write a response that appropriately completes the request. ### Instruction: Complete the following Python code without any tests or explanation [INPUT] ### Response:’.
We evaluated the performance of different watermarks on text quality and watermark strength. For watermark strength, we implemented the $z$-test for all baselines and our methods. We set the $z$ threshold as 2 and 2.5. With $z\geq 2$, we are more than 97.7% confident that the text is watermarked based on the one-tail test.
For text quality, we employed different metrics for different datasets. For the C4 dataset, we utilized perplexity (PPL) and coherence [9] to measure the text quality. For HumanEval, we employed PPL and pass@$k$ score of the code [4]. The pass@$k$ score measures the normalized percentage of solved problems in HumanEval. Formally, the pass score is calculated as
$\text{pass}@k=\mathbb{E}_{\text{Problems}}\big{[}1-\frac{C_{n-c}^{k}}{C_{n}^{k}}\big{]},$
where $c$ is the number of passed codes among $k$ generations.
Baselines. We compared against biased and unbiased watermarks in terms of text quality and watermark strength. For further details of baselines, we refer readers to Appendix A. We implemented all LLMs with the Hugging Face library [32]. All watermark benchmarks including KGW, $\gamma$-reweight, and Dipmark were implemented using their public codes.
Implementation details. For all baselines and our methods, we utilized multinomial sampling during text generation. For C4, we employed LLaMA-2-7B as our generative LLM [29]. Following previous work [12], we continued to sample prompts from C4 until we had generated at least 500 text sequences, each consisting of $T=200\pm 5$ tokens. We leveraged LLaMA-2-13B to compute the perplexity of the generated texts. For HumanEval, we applied CodeLLaMA-7B-Instruct [26] as the generative LLM to generate codes for all Python problems. We also leveraged LLaMA-2-13B to compute the perplexity. All experiments were conducted on a single Nvidia A100 GPU with 80GB memory.
(a) STA-4 on C4
(b) STA-8 on C4
(c) STA-16 on C4
(d) STA-4 on HumanEval
(e) STA-8 on HumanEval
(f) STA-16 on HumanEval
Figure 2: Performance of STA-M w.r.t. $\tau$
Table 5: Examples of STA-generated Texts for C4
Prompt Human-written STA-1 generated STA-16 generated
[]delimited-[][$\dots$][ … ] Single taxpayers who are eligible to participate in a workplace retirement plan are also eligible to make a tax-deductible contribution to an IRA if their adjusted gross income is below \$64,000 (\$103,000 for marrieds) in 2019. This is up from \$63,000 (singles) and \$101,000 (marrieds) in 2018. This deduction is phased out when AGI is between \$64,000 and \$74,000 (singles) and \$103,000 to \$123,000 (marrieds). The income range for making contributions to a Roth IRA in 2019 is \$122,000 to \$137,00 (singles and heads of households) and \$193,000 to \$203,000 (marrieds). The 2019 income limit for the Savers Credit (also called the retirement savings contributions tax credit), which is for low- to middle-income workers who contribute to a retirement plan or IRA, []delimited-[][$\dots$][ … ] (PPL:3.09) and \$74,000 for singles (\$103,000 and \$123,000 for marrieds, respectively). IRA contributions can be made until the 2018 tax-return deadline, April 15, 2018, for those that filed an extension. However, you’ll need to make these contributions with the 2018 deduction in mind. This means you must make IRA contributions no later than Dec. 31, 2018, to benefit on your 2018 return. []delimited-[][$\dots$][ … ] (PPL:3.11) and \$74,000 (\$103,000 and \$123,000 for marrieds) and fully eliminated when AGI exceeds \$74,000 (marrieds phase out at \$123,000). If you’re not able to participate in a 401(k) or other workplace retirement plan, you may qualify to deduct your total IRA contributions even if your income exceeds certain amount if you meet certain conditions (a deductible contributions means you won’t owe tax on the contributions). []delimited-[][$\dots$][ … ]
[]delimited-[][$\dots$][ … ] Thomas will be responsible for overseeing Micron’s solid state storage business that ranges from hard disk drive replacements with solid state drives (SSDs) to enterprise-class storage solutions. He brings more than 30 years of experience to Micron and most recently served as the vice president of Enterprise Storage for Micron’s common stock is traded on the NASDAQ under the MU symbol. To learn more about Micron Technology, Inc., visit www.micron.com. Micron and the Micron orbit logo are trademarks of Micron Technology, Inc. All other trademarks are the property of their respective owners. []delimited-[][$\dots$][ … ] (PPL:3.25) the Americas region for Seagate Technology. He is a senior executive level leader with a proven track record in defining strategy that drives revenue, profit and new technology execution. "Micron is thrilled to have Darren as part of our team," said Mary Jane Raymond, . []delimited-[][$\dots$][ … ] (PPL:4.45) Fusion I/O, LLC. Before that, Thomas was at Western Digital Corporation where he was a progressive executive, holding various management roles since 2008, most recently as its executive vice president of storage technology. []delimited-[][$\dots$][ … ]
[]delimited-[][$\dots$][ … ] Sanabia has benefited from the two times Miami’s offense has given its starters decent run support, including his last outing against Washington. The 24-year-old allowed two runs and six hits over six innings in Tuesday’s 8-2 victory over the Nationals. He tossed six scoreless frames in his only road start against the New York Mets, but is allowing left-handed hitters to bat 8-for-24 against him - a troubling trend against a Reds team that features Choo, Votto and Jay Bruce at the top of the order. []delimited-[][$\dots$][ … ] (PPL:4.30) his prior start at Colorado. Sanoobia is 3-4 with a 4.53 ERA in 13 starts for the Marlins, who are off to the second-worst start in franchise history at 5-13. Johnny Cueto (2-3, 2.63 ERA) was hit around for five earned runs over 6 2/3 innings in a loss to Colorado last Saturday. []delimited-[][$\dots$][ … ] (PPL:5.30) a 5-1 home loss to the L.A. Dodgers eight days earlier. Reds rookie Anthony DeSclafani produced an excellent performance the last time he stepped onto Great American Ball Park. The young right-hander used excellent command of his off-speed pitches to strike out eight []delimited-[][$\dots$][ … ]
Table 6: Examples of STA-generated Texts for HumanEval
Prompt Canonical Solution STA-1 generated STA-16 generated
from typing import List def filter_by_prefix (strings: List[str], prefix: str) -> List[str]: """ Filter an input list of strings only for ones that start with a given prefix. $>$ filter_by_prefix([], ’a’) [] $>$ filter_by_prefix([’abc’, ’bcd’, ’cde’, ’array’], ’a’) [’abc’, ’array’] """ return [x for x in strings if x.startswith(prefix)] (Pass, PPL:1.92) from typing import List def filter_by_prefix (strings: List[str], prefix: str) -> List[str]: """ Filter an input list of strings only for ones that start with a given prefix. $>$ filter_by_prefix([], ’a’) [] $>$ filter_by_prefix([’abc’, ’bcd’, ’cde’, ’array’], ’a’) [’array’, ’abc’] """ return [word for word in strings if word.startswith(prefix)] (Pass, PPL:2.18) from typing import List def filter_by_prefix (strings: List[str], prefix: str) -> List[str]: """* Filter an input list of strings only for ones that start with a given prefix. Input: ‘strings = ["abc", "bcd", "cde", "array"]‘, ‘prefix = ’a’‘ Output: ‘["abc", "array"]‘ """ return [word for word in strings if word.startswith(prefix)]
from typing import List def intersperse (numbers: List[int], delimeter: int) -> List[int]: """ Insert a number ’delimeter’ between every two consecutive elements of input list ‘numbers’ $>$ intersperse([], 4) [] $>$ intersperse([1, 2, 3], 4) [1, 4, 2, 4, 3] """ if not numbers: return [] result = [] for n in numbers[:-1]: result.append(n) result.append(delimeter) result.append(numbers[-1]) return result (Not pass, PPL:2.45) def intersperse (numbers: List[int], delimeter: int) -> List[int]: """ Insert a number ’delimeter’ between every two consecutive elements of input list ‘numbers’ $>$ intersperse([], 4) [] $>$ intersperse([1, 2, 3], 4) [1, 4, 2, 4, 3] """ answer=[] for i in range(0,len(numbers),1): answer = answer + [ numbers[i] ] answer = answer + [ delimeter ] return answer (Not pass, PPL:3.38) Maybe you meant () -> List[int]
def is_prime(n): """Return true if a given number is prime, and false otherwise. ”’ if n < 2: return False for k in range(2, n - 1): if n % k == 0: return False return True (Pass, PPL:1.94) “‘python def is_prime(n): if n == 1: return False for i in range(2, n): if n % i == 0: return False return True (Pass, PPL:1.78) def is_prime(n): if n == 1: return False for i in range(2, n): if n % i == 0: return False return True if __name__ == ’__main__’: import doctest doctest.testmod()
E.2 Robustness Check on Entropy Threshold Parameter
In this section, we conducted a robustness check on the parameter $\tau$ in STA-M. In particular, we set the low entropy threshold $\tau$ from 0 to 2.1 with an interval of 0.15. At each generation step, we sampled at most 4, 8, and 16 times (i.e., STA-4, STA-8, and STA-16) when the entropy was above the threshold $\tau$. Figure 2 shows text quality and watermark strength of STA-M with different $\tau$s. As depicted, different $\tau$s do not affect the watermark strength significantly for C4 because C4 is a high-entropy dataset. Also, we observe a decrease in PPL when we increase $\tau$ in Figure 2(a), 2(b), and 2(c). The reason is that by setting up a higher entropy threshold, fewer generation steps will apply the STA-M strategy, making the watermarking method more similar to STA-1. According to Figure 2(d), 2(e), and 2(f), we observe a general increase of watermark strength if we have a larger $\tau$ because we will have more green list tokens if we sample M times instead of once. However, higher watermark strength leads to a lower pass@$1$ score, which is related to the text quality [12]. We chose the Pareto optimal of each dataset as our final parameter for each dataset. Specifically, we selected $\tau=1.35$ for C4 and $\tau=1.95$ for HumanEval.
E.3 Examples of STA-generated Texts
We present examples of STA-generated texts for C4 and HumanEval in Table 5 and Table 6, respectively. Also, we report the PPL of the generated text, and whether the code is passed specifically for HumanEval.
E.4 Attacking Watermarks
We introduce the implementation of different attacks as follows. For the GPT-3.5 attack, we utilized the prompt ‘Rewrite the following paragraph: [INPUT]’ for GPT-3.5. For DIPPER-1 [14], we set the lexical diversity to 60 without considering order diversity. Additionally, we increased the order diversity by 20 for DIPPER-2 following previous work [17]. For the copy-paste attack, we randomly replaced $25\%$ of tokens in the watermarked text with tokens from non-watermarked text generated from the same prompt [12].
Appendix F Related Work
Existing white-box watermarking techniques fall into two categories: watermarking during logits and probabilities generation, and watermarking by sampling candidate tokens.
Watermarking during logits and probabilities generation. This category of watermarking methods inserts watermarks into LLMs by artificially adjusting the raw logits or probabilities generated by the LLM. Among this category, Kirchenbauer et al. [12] propose the first watermarking method based on logits adjustment. Their approach randomly partitions the vocabulary set into a green and a red list at each generation step, increasing the logits of green list tokens while keeping red list tokens’ logits fixed. Lee et al. [16] extend the green and red list-based watermarking method to low-entropy scenarios. They adjust the logits only during high-entropy generation steps, leaving the raw logits unchanged for low-entropy steps. Ren et al. [25] improve the vocabulary set partition process by determining the green and red lists based on semantic embeddings of preceding tokens rather than their hash values. Fernandez et al. [8] propose a multi-bit watermarking method that generates a multi-dimensional vector at each generation step, which is utilized to modify logits produced by the original LLM. Their approach allows embedding any bit of watermarking information, up to the dimension of the vector used in the logits adjustment. Yoo et al. [34] develop a multi-bit method by extending the two-list partition idea to multi-list partitions. At each generation step, the vocabulary set is divided into multiple lists. Based on the message to be inserted, the logits for tokens in a selected list are increased, while the token logits in all other lists remain unchanged.
Instead of splitting the vocabulary set into different lists, Hu et al. [11] introduce a method that randomly shuffles the order of all token probabilities within the interval $[0,1]$, setting the probabilities in the first half of the interval to 0 and doubling those in the second half. During the detection phase, a likelihood ratio test examines the significance of the likelihood that the given text is generated with the adjusted probability distribution. Wu et al. [33] further generalizes this method by introducing a hyperparameter $\alpha\in[0,0.5]$, which controls the two cutoff points $\alpha$ and $1-\alpha$ within the interval $[0,1]$. The probability masses for the three resulting sub-intervals are adjusted accordingly.
Watermarking by sampling candidate tokens. This category of watermarking methods inserts watermarks into the token sampling process by using watermark information to control the sampling of candidate tokens. For example, Christ et al. [5] introduce a watermarking method that represents each token in the vocabulary set as a binary string of 0s and 1s. Next, a sequence of values from 0 to 1 is sampled uniformly. These values guide the token sampling process: if the predicted probability for a position in the binary string is larger than the corresponding pseudo-random value, that position is assigned a 1; otherwise, it is assigned a 0. Once all positions are determined, the token corresponding to the resulting binary string is sampled. Additionally, previous work [15] use a sequence of values randomly sampled from a uniform distribution between 0 and 1. The value controls the token sampling process through a decoder function, where the decoder function varies based on the sampling strategy. Hou et al. [10] sample new sentences according to the original LLM until a sentence’s semantic value falls into the acceptance region. The acceptance region is predefined by randomly splitting the space of semantic embedding according to the context and the key.